Calculate the pH of a 0.0035 M solution of HNO3.
1
Expert's answer
2020-05-01T14:42:10-0400

[OH-] = 0.0035 M

pH = 14 - pOH

pOH = -Log[OH-] = 2.456

pH = 14 - 2.456 = 11.544